OpenStudy (anonymous):

Find the limit of sqrt{x^2 - 1}/ x

OpenStudy (anonymous):

As x approaches infinity

OpenStudy (anonymous):

that would be 1. what are you allowed to use?

OpenStudy (anonymous):

probably simplest to say it is \[\lim_{x\rightarrow \infty}\sqrt{\frac{x^2-1}{x^2}}\] which is pretty clearly 1 since you can take the limit inside the square roots